I have an Epidemiology question, how do you calculate prevalence with this data total population 8,630,000? And do you change the quantity of the population when there is deaths by subtracting? Do I use total case/population? Thank you!

Date    Total case New case    Total death New death

1/17/2015 462 231    9    5

3/30/2015    1386 924    28    18
Answer #1

Answer :

Prevalance formula in Epidemiology :

Number of individuals identified as cases during in the specified period of time / total number of people in that population.

Here

* Total population 86,30,000.

* total deaths we will subtract here.

* on 1 / 17/ 2015 , total deaths are 9 + 5= 14.

* if we subtract from total population 86,30,000 - 14 = 86,29,986.

* Total cases are ( new and old ) = 693.

* Prevalance = 693 / 86,29,986 = 0.0000803014.

* prevalence rate means we need to multiply with 100.

* so 0.008030.

* On 3 / 30 /2015 , total Deaths are 28 + 18= 46.

* if we subtract from total population 86,30,000 - 46 = 86,29,954.

* Total cases ( old and new in that period)= 1386 + 924 = 2310.

* Prevalance = 2310/ 8629954 = 0.000267672 .

* Prevalance rate means we need multiply with 100.

* so 0.02676.